OpenStudy (anonymous):

5x^2+16x+9=0 solve by factoring

OpenStudy (anonymous):

\[if a \not = 0 , ax^2+bx+c = 0\] then we solve for x in this way: \[\Delta = b^2-4ac\] \[x = (-b \pm \sqrt{\Delta})/2a\]

OpenStudy (anonymous):

im confused